### v2

@@ Abstract
This paper outlines the existing situation and future trends related to automobile battery packs, specifically from the automobile manufacturer’s point of view. It formulates the specifications required for such packs to adhere to prevailing regulatory schemes (ISO 26262, UN ECE R100) and examines top-level solutions to target a uniform architecture for passenger cars. Key elements such as electrical performance, safety, mechanical integrity, reliability, environmental issues, diagnostics, and real-world implications have been extensively examined. This paper draws attention to the industry trend of shifting to high-voltage battery architectures (400V–800V) to enable ultra-fast charging above 350 kW, reducing the charging time to less than 20 minutes. Technological advancements in energy density (up to 270 Wh/kg) and battery pack capacities (100–150 kWh) are poised to take electric vehicle ranges over 1000 km from a single charge. This study also examines developments in artificial intelligence-improved battery management systems, enhanced safety, mechanical integrity, reliability, diagnostics, and practical considerations. Furthermore, future developments, such as the incorporation of batteries in aviation and other new uses, are investigated to provide insight into the future generation of economically viable, secure, and high-performance battery systems.
@@
Battery pack Automotive Status review Technological trends
@@ 1 Article Highlights
  • Providing the latest technological trends and near future developments for automotive battery pack
@@ 2 Introduction
Battery packs are key components of electric vehicles (EVs) because they operate as the main power supply. Despite recent advancements, further improvements are required to achieve smaller, cheaper, and safer units. Fig. 1 shows the ideal battery pack and major constraints. The battery pack, as the main energy storage device for EVs, delivers the required energy and power with a reliable and durable operation that is safe and environmentally friendly. In addition, fast charging is a highly required feature by customers, which adds new aspects to battery pack design, such as busbar temperature monitoring. Different studies have discussed various aspects of batteries Matthew ; Feng ; Naguib ; Xie ; Hu ; Teichert ; Ghaeminezhad ; Kuang ; Ruffa ; Roscher . However, there is still no comprehensive public report showing the overall requirements and development process in the automotive industry. The gap between the automotive industry and academia is well known. This work aims to provide a comprehensive and compact overview and status review of battery packs and near-future trends from an automotive perspective. However, the focus is on the required specifications and regulatory standards, which are driving forces for available solutions. This review aims to bridge the gap between academic research and industry requirements by providing a structured analysis of automotive battery pack standards, key design aspects, and emerging trends. This article is intended to serve as a reference for researchers, engineers, and policymakers in the field of EV battery technology. Additionally, it aims to highlight the latest advancements and identify future challenges that need to be addressed for the development of safer, more efficient, and cost-effective battery systems.
@@
Currently, electrification is the subject of heavy investment in the automotive industry. Many activities and developments are underway, but there are limited publicly available data. The main goal of this work is to provide an automotive side view of the topic with a focus on applicable standards from the vehicle level down to the component level. In addition, recent developments and trends are provided at an appropriate level that is suitable in this context. This paper is organized into two major parts: different requirements and available solutions that consider regulatory standards. Future trends have been added to each relevant section to ensure greater consistency. Each part is divided into several sections that cover the vehicle level to the component level if required. The main task of battery storage is to provide energy and power in predefined voltage and current windows. These electrical requirements are explained in Section II. However, there are other specifications such as the cooling method, charging and discharging capabilities, power quality, electromagnetic compatibility (EMC), and more similar requirements that are addressed in this section. The next section is dedicated to the battery pack interfaces toward the vehicle. Power and signal connectors and cooling interfaces are the subjects of this section. The final battery pack should be able to operate in a harsh automotive environment, which is mainly governed by ISO 16750 ISO16750-1 ; ISO16750-2 ; ISO16750-3 ; ISO16750-4 ; ISO16750-5 . The battery pack was subjected to extensive environmental testing, such as temperature, vibration, and humidity. This is discussed in Section IV. Safety is one of the most important requirements of automotive battery packs, as discussed in Section V. The battery pack should be electrically and mechanically safe, and different criteria should be fulfilled as required by the standards. Functional safety is also the main tool for realizing the requirements mentioned. In response to all the desired requirements, an electrical architecture should be designed and implemented for the battery pack. The battery system architecture for automotive applications converges to a classical structure; however, there have been some slight changes. The battery system structure is discussed in Section VI. The cell technology, battery modules, and power stage of the battery system are presented in this section. A battery management system (BMS) is a battery system controller that performs measurements, supervision, control, and communication. Section VII is describes the BMS and its issues. Some new aspects, such as cybersecurity requirements, were also covered. Section VIII discusses new applications and general topics. In addition, future studies are presented. The final section presents the conclusions.
@@ 3 Electrical Requirements
The electrical specifications of the battery pack as the source of traction energy and power are explained in this section. The battery pack should be able to provide the required power and energy for a predetermined lifetime or operational cycle. In addition, the capacity must be within the specified values during different operating conditions and should remain stable during the lifecycle with limited aging factors. All these features are part of the requirements discussed in this section.
@@ 3.1 Voltage, Energy and Power
A nominal battery voltage of 350 V, 400 V, 700 V, or even higher is used worldwide for EVs, depending on the vehicle 400_battery . However, the traction force demand is constantly increasing for different commercial and technical reasons. The price of components reduces over time, and performance is constantly improving in terms of weight and efficiency. Battery systems with higher voltages, that is, 700 V are expected to become more popular in upcoming years. The vehicle electrical system can be assumed to be a small power generation and distribution system, with similar issues. Increasing the system voltage can reduce losses for a well-designed system. More importantly, with a higher voltage level, it is easier to perform fast charging.
The battery voltage is subjected to a considerable change from the state of being fully charged to the state of depletion. Consequently, nominal values are used for system identification and are not typically used in the component design stage. For example, the minimum and maximum voltage values were used for the design of an electric motor. However, increasing the system voltage is not an easy task. There are different aspects, such as safety, that limit the components, both technical and commercial. Hence, it is important to monitor the standard regulations for voltage levels in the automotive industry. ISO 6469 is a widely used safety standard for energy storage systems that has four parts ISO6469-1 ,ISO6469-2 ,ISO6469-3 , and ISO6469-4 . There are two main classes of voltage for automotive applications; classes A and B. However, Class B had more subclasses. Table 1 lists automotive voltage classes ISO6469-3 . As shown in this table, under 60 V of DC is class A, and over 60 V DC to 1500 V is class B. Hence, the battery pack for traction applications is mostly class B. There are different safety requirements for the B1 and B2 classes. Class B2 is targeting 48 V systems.
@@ B $60\leq U\leq 1500$ $30\leq U\leq 1000$
In the automotive case, the system status is known for all the cases. In addition, the system’s safety status or functional actions should be predetermined and implemented. Hence, the required system performance under different voltage conditions should be clarified, as shown in Fig. 2 as a typical example. Some areas in the figure may have been integrated. For example, high voltage and overvoltage may occur in the same regions. As a practical example, for a system with a 350 V nominal value, the nominal range can be 300-420 V which the system performs fully in terms of torque and power. The vehicle should be able to operate at low voltages up to 200 V but with limited torque. For high voltages, it can be expected to have a system with reduced performance up to 480 V and afterward, a hard protection action is needed above this voltage level. Battery energy and power are important electrical requirements that should be specified for the vehicle lifecycle. It is well known that cells are subject to aging, and battery capabilities diminish over the operating cycle. The energy and power of the battery were specified at the beginning and end of its life. There might even be more terms, such as the middle of life in some companies. Another important factor is the impact of temperature. Available energy is significantly reduced under cold weather conditions. Additionally, high temperatures pose safety risks. Around room temperatures (15-35 ${}^{o}C$) available power and energy are full, while they shall be derated at low and high temperatures Karlsen ; Lajunen1 ; Lajunen2 ; Jones ; Sundin . A maximum power level of 100-200 kW can be assumed for a passenger EV. Moreover, several ongoing developments are aimed at reaching a 1000 km per charge. It is expected that an energy level of 100-150 kWh can be seen in many upcoming vehicles. There are different standards covering the battery performance of the IEC 62660 series ISO62660-1 ; ISO62660-2 ; ISO62660-3 , ISO 12405 series ISO3 ; ISO4 , UL 2580 UL2580 , SAE J 2929 SAEJ2929 , and GB/Z 1833.1 GBZ1833 are examples. Further details are provided below. The powertrain unit of an EV can be a front-wheel drive, rear-wheel drive, or four-wheel drive. In front- or rear-wheel drive there, is a single electric motor and inverter, whereas in four-wheel drive, there are usually two electric motors and inverters. Hence, there can be different power demand requirements for the batteries. From a vehicle perspective, each platform is usually designed for three power levels: low, medium, and high duty. For a passenger car, the values for low, medium, and high duties can be 80 kW, 100 kW, and 120 kW peak power, respectively. These numbers were approximate. However, in sports vehicles and new designs, there is a trend for higher power levels. For example, Polestar vehicles may reach a peak power of 500 kW in all-wheel drive configurations. The increasing powertrain demand towards higher power levels and easier fast charging are the main motivations for moving towards higher voltage levels like 800 V. However, there is an optimal compromise for component losses and practical safety considerations that limit the maximum voltage.
@@ 3.2 Charging, discharging, and life cycle
The charging and discharging requirements of the battery pack are directly related to the power demand required by the electric motors and the charging time. The battery pack design shall be such that could meet the required max power in traction and regeneration modes. In addition, the charging power is a critical factor for end users. A passenger vehicle typically has a lifetime of approximately 15 years or can travel up to 300,000 kmdistance ,Americans . The idea is to design a battery pack to fulfill these requirements. For EVs, based on the end-user usage behavior statistics, these numbers are somehow converted to the number of deep charge and discharge cycles or hours of operation for the electronics. The battery capacity should be sized based on the target mission profile, depending on the type of vehicle which can be hybrid or electric. The electronics connected to the battery or other components have a lifetime of 40,000-100,000 h. The capacity of a cell is determined based on available statistics, including driver behavior. However, in the specification, the capacity will be specified for extreme conditions, such as low-temperature regions. Chargers, onboard or offboard, are critical components for the market penetration of EVs. Hence, they have been the subject of extensive research and development Artex . Refer to related articles for more details and standards. There are different standards for charging, including auxiliary systemssuch as signaling and earth monitoring, which can be found in references IEC61851_1 ; IEC61851_21 ; IEC61851_22 ; IEC61851_23 ; IEC61851_24 ; IEC62196_1 ; IEC62196_2 ; IEC62196_3 ; SAEJ1772 ; SAEJ2293_1 ; SAEJ2293_2 ; UL2231_1 ; UL2231_2 ; UL2251 ; UL2202 ; GBT18487_1 ; GTB1847_2 ; GTB1847_3 ; CHAdeMO . Chargers are classified as either conductive or wireless. Wireless chargers are under development, but they still require further improvement in terms of cost and weight. Hence, they are not popular. Conductive chargers are the dominant solution. However, some attempts have been made to introduce other forms of charging, such as charging via roads, to reduce the required battery capacity. The onboard chargers are in the power range of 3-22 kW. Because of its cost and weight, it is difficult to reach higher power levels. However, with new power switches, such as SiC devices, there is a good chance to see higher power levels soon. Offboard chargers are commercially available in the MW range for many companies. For passenger cars, the trend is towards 350 kW chargers. However, currently, most vehicles are not ready to absorb this power level. There are power limits exist for the battery pack, connectors, and connections. The subject is under extensive development by automakers, and a power level of 200 kW is currently available for several vehicles. Another class of chargers is integrated motor drive and battery chargers. This technology provides higher power charging available onboard, but has some practical difficulties for implementation. Because they are non-isolated, safety and power quality issues are challenging Haghbin1 . However, there are some commercially available products like ZOE that, use integrated motor drives and battery chargers. This alternative has a good chance of being widely used shortly because it offers an onboard fast charger with a low price and compact volume. The battery pack should be able to operate under predetermined conditions with specified charging and discharging levels in all temperature and environmental conditions. Usually, the lifetime is determined for as the beginning of the lifetime, the middle of the lifetime, and the end of the lifetime. In addition, based on the available internal data, there are company-specific requirements in line with the standard requirements.
@@ 3.4 EMC
The battery pack, as an individual component with connectors and interfaces, including all cells and electronics, has an acceptable EMC behavior, as defined in relevant standards. The different types of emissions were limited to a certain level for different frequency ranges. Since the battery pack has a metallic enclosure, it is usually critical to have good connectors for EMC behavior. There may be some shielding and the addition of an XY capacitor to enhance this behavior. EMC filters are very common in different parts of the system. For example, it is common to have an EMC filter in the DC output of a pack. The pack radiation can be measured practically using the recommended setup, as stated in the standard. However, the main challenge is to correlate EMC behavior with the components. For example, if there is a peak in radiation at a certain frequency, it is challenging to find the source of this pollution. Because of the complex nature of the EMC at the pack level, practical measurement and trial and error are widely used to enhance EMC performance. Better analytical or simulation models are required to predict the system EMC behavior more accurately. There are many standards that the vehicle should fulfill, depending on the target country ISO11451_1 ; ISO11451_2 ; ISO11451_3 ; ISO11451_4 ; ISO11452_1 ; ISO11452_2 ; ISO11452_3 ; ISO11452_4 ; ISO11452_5 ; ISO11452_7 ; ISO11452_8 ; ISO11452_9 ; ISO11452_10 ; ISO11452_11 ; ISO7637_1 ; ISO7637_2 ; ISO7637_3 ; ISO10605 ; CISPR12 ; CISPR25 ; CISPR36 . The standards describe the intended behavior in different frequency ranges. However, automakers may have internal requirements that are usually more stringent than standard requirements. Usually, pack-level EMC is handled by the automaker, but at the component level, the supplier will have the main responsibility. For example, the BMS supplier should deliver the part with predetermined EMC levels or standard requirements. The Automaker can then judge at the pack level how to change the limit or manage the EMC at the pack level. This approach requires enhancements to achieve targets. Occasionally, a practical trial and error is insufficient for the pack level. More research is needed on this subject. Another point worth mentioning is that the entire vehicle was subjected to the EMC test. The results were then correlated to the subsystem level. It is important to analyze the spectrum to determine the source of violations. Afterward, a mitigation strategy is prepared and implemented.

### v3

@@ Abstract
The latest advancements and near-future trends in automotive battery packs, underlying regulatory compliance, and performance requirements are presented in this paper. In response to these specifications, high-level solutions that converge towards a standard architecture for passenger cars are provided. Transition to high-voltage (400V–800V) enables ultra-fast charging above 350 kW, which reduces the charging times to less than 20 minutes. Also, advances in energy density (up to 300 Wh/kg) and battery capacities make advancements in enhancing the electric vehicle’s range beyond 1000 km per charge. Key factors such as electrical performance, safety, mechanical integrity, reliability, endurance, environmental conditions, and diagnostics are examined. This study explores the next generation of cost-effective and high-performance battery systems and discovers near-future battery technologies, including sodium-ion chemistry and rare-earth-free alternatives, as well as battery applications in aviation.
@@
Automotive Battery pack Electric Vehicle Status review Technological trends
@@ 1 Article Highlights
  • Providing the latest technological trends and near-future developments for automotive battery pack
@@ 2 Introduction
Battery packs are key components of electric vehicles (EVs) because they operate as the main power supply. Despite recent advancements, further improvements are required to achieve smaller, cheaper, and safer units Matthew ; Feng ; Naguib . Fig. 1 shows the ideal battery pack and major constraints. The battery pack, as the main energy storage device for EVs, delivers the required energy and power with a reliable and durable operation that is safe and environmentally friendly Xie ; Hu . In addition, fast charging is a highly required feature by customers, which adds new aspects to battery pack design, such as busbar temperature monitoring. Different studies have discussed various aspects of batteries Teichert ; Ghaeminezhad ; Kuang ; Ruffa ; Roscher . However, there is still no comprehensive public report showing the overall requirements and development process in the automotive industry. The gap between the automotive industry and academia is well known. This work aims to provide a comprehensive and concise overview and status review of battery packs and near-future trends in the automotive industry. In particular, the required specifications and regulatory standards are more interested. This review seeks to connect academic research with industry needs by offering a comprehensive overview of automotive battery pack standards, developments in that field, and the essential design considerations. It can serve as a reference for researchers and policymakers in the EV battery technology, which aims to provide more information the latest advancements and identify future challenges that need to be addressed for the development of safer, more efficient, and cost-effective battery systems.
@@
Currently, electrification is the subject of heavy investment in the automotive industry. Many activities and developments are underway, but there are limited publicly available data. The main goal of this work is to provide an automotive side view of the topic with a focus on applicable standards from the vehicle level down to the component level. In addition, recent developments and trends are provided at an appropriate level that is suitable in this context. This paper is organized into two major parts: different requirements and available solutions that consider regulatory standards. Future trends have been added to each relevant section to ensure greater consistency. Each part is divided into several sections that cover the vehicle level to the component level if required. The main task of battery storage is to provide energy and power in predefined voltage and current windows. These electrical requirements are explained in Section II. However, there are other specifications, such as the cooling method, charging and discharging capabilities, power quality, electromagnetic compatibility (EMC), and more similar requirements that are addressed in this section. The following section is dedicated to the battery pack interface with the vehicle. Power and signal connectors and cooling interfaces are the subjects of this section. The final battery pack should be able to operate in a harsh automotive environment, which is mainly governed by ISO 16750-1 ISO16750-1 , ISO16750-2 ISO16750-2 , ISO ISO16750-3 ISO16750-3 , ISO 16750-4 ISO16750-4 , and ISO 16750-5 ISO16750-5 , offering guidance on typical environmental factors that affect electrical and electronic systems in automobiles, along with specified requirements and testing procedures. The battery pack was subjected to extensive environmental testing, such as temperature, vibration, and humidity. This is discussed in Section IV. Safety is one of the most important requirements of automotive battery packs, as discussed in Section V. The battery pack should be electrically and mechanically safe, and different criteria should be fulfilled as required by the standards. Functional safety is also the main tool for realizing the requirements mentioned. In response to all the desired requirements, an electrical architecture should be designed and implemented for the battery pack. The battery system architecture for automotive applications converges to a classical structure; however, there have been some slight changes. The battery system structure is discussed in Section VI. The cell technology, battery modules, and power stage of the battery system are presented in this section. A battery management system (BMS) is a battery system controller that performs measurements, supervision, control, and communication. Section VII describes the BMS and its issues. Some new aspects, such as cybersecurity requirements, were also covered. Section VIII discusses new applications and general topics. In addition, future studies are presented. The final section presents the conclusion.
Figure 2: 3D view of a battery cell and a battery pack used in an electric vehicle battery_pack_2024 .
Figure 3: Tesla 5.32 kWh battery module (85 kWh battery pack).
@@ 3 Electrical Requirements
Fig. 2 shows 3D view of a battery cell and battery pack which can be used in an EV. Fig. 2 shows a 3D view of a battery pack in an EV. Fig. 3 shows the Tesla battery module with a capacity of 5.3 kWh, which contributes to a battery pack of 85 kWh. Electrical specifications of the battery pack as the source of traction energy and power are explained in this section. The battery pack should be able to provide the required power and energy for a predetermined lifetime or operational cycle. In addition, the capacity must be within the specified values during different operating conditions and should remain stable during the lifecycle with limited aging factors. All these features are part of the requirements discussed in this section. Electrical characteristics of a battery pack reveal its ability to deliver consistent power and energy throughout its lifespan. The battery system should be stable under different conditions, and consider the minimization of the battery pack aging effects to preserve performance and reliability. These factors play a key role in the efficient and reliable design of the battery systems of EVs.
@@ 3.1 Voltage, Energy and Power
A nominal battery voltage of 350 V, 400 V, 700 V, or even higher is used worldwide for EVs, depending on the vehicle 400_battery . However, the traction force demand is constantly increasing for different commercial and technical reasons. The price of components reduces over time, and performance is constantly improving in terms of weight and efficiency. Battery systems with higher voltages, that is, 700 V, are expected to become more popular in upcoming years. The vehicle electrical system can be assumed to be a small power generation and distribution system, with similar issues. Increasing the system voltage can reduce losses for a well-designed system. More importantly, with a higher voltage level, it is easier to perform fast charging.
The battery voltage is subjected to a considerable change from the state of being fully charged to the state of depletion. Consequently, nominal values are used for system identification and are not typically used in the component design stage. For example, the minimum and maximum voltage values were used for the design of an electric motor. However, increasing the system voltage is not an easy task. There are different aspects, such as safety, that limit the components, both technical and commercial. Hence, it is important to monitor the standard regulations for voltage levels in the automotive industry. ISO 6469 is a widely used safety standard for energy storage systems and protection against failures caused by hazards unique to electrically propelled road vehicle that has four parts ISO6469-1 ISO6469-1 , ISO 6469-2 ISO6469-2 , ISO 6469-3 ISO6469-3 , and ISO 6469-4 ISO6469-4 . There are two main classes of voltages for automotive applications; classes A and B. However, Class B had more subclasses. Table 1 lists automotive voltage classes ISO6469-3 . As shown in this table, under 60 V of DC is class A, and over 60 V DC to 1500 V is class B. Hence, the battery pack for traction applications is mostly class B. There are different safety requirements for the B1 and B2 classes. Class B2 is targeting 48 V systems.
@@ B $60\leq U\leq 1500$ $30\leq U\leq 1000$
In the automotive case, the system status is known for all the cases. In addition, the system’s safety status or functional actions should be predetermined and implemented. Hence, the required system performance under different voltage conditions should be clarified, as shown in Fig. 4 as a typical example. Some areas in the figure may have been integrated. For example, high voltage and overvoltage may occur in the same regions. As a practical example, for a system with a 350 V nominal value, the nominal range can be 300-420 V which the system performs fully in terms of torque and power. The vehicle should be able to operate at low voltages up to 200 V but with limited torque. For high voltages, it can be expected to have a system with reduced performance up to 480 V and afterward, a hard protection action is needed above this voltage level. Battery energy and power are important electrical requirements that should be specified for the vehicle lifecycle. It is well known that cells are subject to aging, and battery capabilities diminish over the operating cycle. The energy and power of the battery were specified at the beginning and end of its life. There might even be more terms, such as the middle of life, in some companies. Another important factor is the impact of temperature. Available energy is significantly reduced under cold weather conditions. Additionally, high temperatures pose safety risks. Around room temperatures (15-35 ${}^{o}C$), available power and energy are full, while they shall be derated at low and high temperatures Karlsen ; Lajunen1 ; Lajunen2 ; Jones ; Sundin . A maximum power level of 100-200 kW can be assumed for a passenger EV. Moreover, several ongoing developments are aimed at reaching a 1000 km per charge. It is expected that an energy level of 100-150 kWh can be seen in many upcoming vehicles. Various standards regulate battery performance, such as the IEC 62660 series, which encourages manufacturers to design interchangeable batteries based on a standardized format, including ISO 62660-1 ISO62660-1 , ISO62660-2 ISO62660-2 , ISO62660-3 ISO62660-3 ; ISO 12405 which is designed for the approval of EVs, including ISO 12450-3 ISO3 , and ISO 12450-4 ISO4 ; UL 2580, which is testing Standard for batteries of EVs, UL2580 ; SAE J2929, which establishes the minimum safety requirements for lithium-ion battery systems used in vehicle propulsion applications connected to a high-voltage powertrain SAEJ2929 ; and GB/Z 1833.1 GBZ1833 . Further details are provided below. The powertrain unit of an EV can be a front-wheel drive, rear-wheel drive, or four-wheel drive. In front- or rear-wheel drive there is a single electric motor and inverter, whereas in four-wheel drive, there are usually two electric motors and inverters. Hence, there can be different power demand requirements for the batteries. From a vehicle perspective, each platform is usually designed for three power levels: low, medium, and high duty. For a passenger car, the values for low, medium, and high duties can be 80 kW, 100 kW, and 120 kW peak power, respectively. These numbers were approximate. However, in sports vehicles and new designs, there is a trend for higher power levels. For example, Polestar vehicles may reach a peak power of 500 kW in all-wheel drive configurations. The increasing powertrain demand towards higher power levels and easier fast charging are the main motivations for moving towards higher voltage levels like 800 V. However, there is an optimal compromise for component losses and practical safety considerations that limit the maximum voltage.
@@ 3.2 Charging, discharging, and life cycle
The charging and discharging requirements of the battery pack are directly related to the power demand by the electric motors and the charging time. The battery pack design shall be such that it could meet the required maximum power in traction and regeneration modes. In addition, the charging power is a critical factor for the end users. A passenger vehicle typically has a lifetime of approximately 15 years or can travel up to 300,000 km distance ,Americans . The idea is to design a battery pack to fulfill these requirements. For EVs, based on the end-user usage behavior statistics, these numbers are somehow converted to the number of deep charge and discharge cycles or hours of operation for the electronics. The battery capacity should be sized based on the target mission profile, depending on the type of vehicle, which can be hybrid or electric. The electronics connected to the battery or other components have a lifetime of 40,000-100,000 h. The capacity of a cell is determined based on available statistics, including driver behavior. However, in the specification, the capacity will be specified for extreme conditions, such as low-temperature regions. Chargers, onboard or offboard, are critical components for the market penetration of EVs. Hence, they have been the subject of extensive research and development Artex . Refer to related articles for more details and standards. There are different standards for charging, including auxiliary systemssuch as signaling and earth monitoring, which can be found in references IEC61851_1 ; IEC61851_21 ; IEC61851_22 ; IEC61851_23 ; IEC61851_24 ; IEC62196_1 ; IEC62196_2 ; IEC62196_3 ; SAEJ1772 ; SAEJ2293_1 ; SAEJ2293_2 ; UL2231_1 ; UL2231_2 ; UL2251 ; UL2202 ; GBT18487_1 ; GTB1847_2 ; GTB1847_3 ; CHAdeMO . Chargers are classified as either conductive or wireless. Wireless chargers are under development, but they still require further improvement in terms of cost and weight. Hence, they are not popular. Conductive chargers are the dominant solution. However, some attempts have been made to introduce other forms of charging, such as charging via roads, to reduce the required battery capacity. The onboard chargers are in the power range of 3-22 kW. Because of its cost and weight, it is difficult to reach higher power levels. However, with new power switches, such as SiC devices, there is a good chance to see higher power levels soon. Offboard chargers are commercially available in the MW range for many companies. For passenger cars, the trend is towards 350 kW chargers. However, currently, most vehicles are not ready to absorb this power level. There are power limits exist for the battery pack, connectors, and connections. The subject is under extensive development by automakers, and a power level of 200 kW is currently available for several vehicles. Another class of chargers is integrated motor drive and battery chargers. This technology provides higher power charging available onboard, but has some practical difficulties for implementation. Because they are non-isolated, safety and power quality issues are challenging Haghbin1 . However, there are some commercially available products like ZOE that, use integrated motor drives and battery chargers. This alternative has a good chance of being widely used shortly because it offers an onboard fast charger with a low price and compact volume. The battery pack should be able to operate under predetermined conditions with specified charging and discharging levels in all temperature and environmental conditions. Usually, the lifetime is determined for as the beginning of the lifetime, the middle of the lifetime, and the end of the lifetime. In addition, based on the available internal data, there are company-specific requirements in line with the standard requirements.
@@ 3.4 EMC
The battery pack, as an individual component with connectors and interfaces, including all cells and electronics, has acceptable EMC behavior, as defined in relevant standards. The different types of emissions are limited to a certain level for different frequency ranges. Since the battery pack has a metallic enclosure, it is usually critical to have good connectors for EMC behavior. There may be some shielding and the addition of an XY capacitor to enhance this behavior. EMC filters are commonly used in different parts of the system. For example, it is common to have an EMC filter in the DC output of a pack. However, the most significant source of electromagnetic interference (EMI) in EV, particularly concerning homologation challenges, is the on-board charger. Due to its high switching frequencies and power levels, the on-board charger generates both conducted and radiated emissions that can interfere with other vehicle systems. As a result, special attention must be given to its EMC design, including proper shielding, filtering, and compliance validation. To ensure adherence to UN ECE Regulation No. 10, which governs the electromagnetic compatibility of road vehicles, the emissions and immunity performance of the on-board charger and other key vehicle components must be tested against standardized limits. This regulation specifies acceptable EMI levels and testing procedures to ensure the safe and interference-free operation of electrical systems.
Practical EMC testing conditions follow standardized methodologies, such as those outlined in ISO 11451, serves as a foundation for program owners, regulators, and accreditation bodies to evaluate and acknowledge the competence of validation and verification organizations, including ISO 11451-1 ISO11451_1 , ISO 11451-2 ISO11451_2 , ISO 11451-3 ISO11451_3 , ISO 11451-4 ISO11451_4 , ISO 11452, which is a collection of international standards and guidelines for testing the immunity of automotive electrical components to narrowband radiated electromagnetic interference from external sources, including ISO 11452-1 ISO11452_1 , ISO 11452-2 ISO11452_2 , ISO 11452-3 ISO11452_3 , ISO 11452-4 ISO11452_4 , ISO 11452-5 ISO11452_5 , ISO 11452-7 ISO11452_7 , ISO 11452-8 ISO11452_8 , ISO 11452-9 ISO11452_9 , ISO 11452-10 ISO11452_10 , and ISO 11452-11 ISO11452_11 , ISO 7637, which is relates to 12- and 24-volt electrical systems, consisting of ISO 7637-1 ISO7637_1 , ISO 7637-2 ISO7637_2 , ISO 7637-3 ISO7637_3 , and ISO 10605 ISO10605 , and CISPR (CISPR12 CISPR12 , CISPR25 CISPR25 , and CISPR36 CISPR36 ), which define measurement techniques for both conducted and radiated emissions, and related regulating electromagnetic interference in electrical and electronic devices as part of the International Electrotechnical Commission. These tests utilize shielded enclosures, bulk current injection, and reverberation chambers to assess compliance and more effectively pinpoint sources of EMI.
@@
Pack radiation can be measured practically using the recommended setup, as stated in the standard. However, the main challenge is correlating EMC behavior with individual components. For example, if there is a peak in radiation at a certain frequency, it is challenging to find the exact source of interference due to the complex nature of EMC at the pack level, practical measurement and trial and error are widely used to enhance EMC performance. There are many standards that the vehicle should fulfill, which depend on each country. The standards describe the intended behavior in different frequency ranges. However, automakers may have their internal requirements that are usually tougher than other standard requirements. Usually, pack-level EMC is handled by the automaker, but at the component level, the supplier will have the main responsibility. For example, the BMS supplier should deliver the part with predetermined EMC levels or standard requirements. The automaker can then assess the EMC at the pack level and decide whether additional mitigation strategies are needed. Occasionally, practical trial and error is insufficient at the pack level, so more research is required on this subject. In addition, the vehicle is examined for EMC testing, and then, the results are correlated to the subsystem level. The spectrum analysis is important since the detection of the violation’s source is concerned. Afterward, a mitigation strategy is prepared and implemented.